OpenStudy (anonymous):

x+9/5x^2+6x+1 plus x/5x^2+7x+2

OpenStudy (anonymous):

\[\frac{x+9}{5x^2+6x+1}+\frac{x}{5x^2+7x+2}\] \[=\frac{x+9}{(x+1)(5x+1)}+\frac{x}{(x+1)(5x+2)}\]

OpenStudy (anonymous):

ok i lied, there is a common factor. it is x+1

OpenStudy (anonymous):

so least common denominator will be \[(x+1)(5x+1)(5x+2)\]

OpenStudy (anonymous):

mulitply top and bottom of first fraction by 5x+2 and second by 5x+1

OpenStudy (anonymous):

\[\frac{(x+9)(5x+2)}{(x+1)(5x+1)(5x+2)}+\frac{x(5x+1)}{(x+1)(5x+1)(5x+2)}\]

OpenStudy (anonymous):

=\[\frac{(x+9)(5x+2)+x(5x+1)}{(x+1)(5x+1)(5x+2)}\]

OpenStudy (anonymous):

now mulitply out in the numerator and combine like terms you get for the numerator \[10x^2+48x+18\]

OpenStudy (anonymous):

so your 'final answer" is \[\frac{10x^2+48x+18}{(x+1)(5x+1)(5x+2)}\]

OpenStudy (anonymous):

No! x+9/5x^2+6x+1 plus x/5x^2+7x+2 means x+(9/5x^2)+6x+1 + (x/5x^2)+7x+2

OpenStudy (anonymous):

i doubt it

OpenStudy (anonymous):

It's what the OP wrote. If OP meant it differently, then parens needed to be used.

OpenStudy (anonymous):

that is why there were common factors in the denominator, so you can realize that the common denominator need not be the product of the two denominators. just take the factors

OpenStudy (anonymous):

You can't assume that. That's what parens are meant for. lol